Which of the following DNA samples has the highest melting temperature?

DNA that is 40% C
DNA that is 30% T
DNA that is 30%G
DNA that is 40%A
Answer #1

a) DNA that is 40% C

According to Chargraff`s rule

(A) +(G) =(C)+(T)

(A) +(T)+(C)+(G) = 100%

As cytosine (C) pairs with Gaunine (G), therefore the proportional of gaunine should be equal to cytosine i,e

(C) =40%, (G) =40%,

Total GC content is 80%, and remaing 20% is AT content, it will have highest melting temperature by putting these values in following formula:

Tm = ( 4 x [G+C] ) + (2 x [A+T] ) 0C